# Difference-in-Discontinuities: Estimation, Inference and Validity Tests

### v1

@@ 4.2 Testing the Time-Invariance of Data-Generating Functional Form
Therefore, it can be very beneficial for researchers to know if the shapes of the data-generating functions remain stable over time. This knowledge would allow them to determine if they can use simpler and less biased estimation methods. To test this, we propose comparing the derivatives of the estimated functions on each side of the threshold. We can estimate $\mu_{+,t}(z)$ and $\mu_{-,t}(z)$ nonparametrically, for $t=\{0,1\}$. This is, we locally approximate $\mu_{t}(z)$ by a polynomial of order $p$ for z in the neighborhood of $\bar{z}$:
$\mu_{t}(z)\sim\mu_{t}(\bar{z})+\mu^{(1)}_{t}(\bar{z})\left(z-\bar{z}\right)+...+\frac{\mu^{(p)}_{t}(\bar{z})}{p!}\left(z-\bar{z}\right)^{p}$
We perform a local polynomial regression fit for each side of the threshold and each period $t={0,1}$:
$\hat{\beta}_{+,t}$ $=arg\min_{\beta\in\mathbf{R}^{p+1}}\sum_{i=1}^{n}\mathbbm{1}(Z_{i}\geq 0)(Y_{i,t}-r_{p}(Z_{i})^{\prime}\beta)^{2}K_{h_{n}}(Z_{i}-\bar{z})$
$\hat{\beta}_{-,t}$ $=arg\min_{\beta\in\mathbf{R}^{p+1}}\sum_{i=1}^{n}\mathbbm{1}(Z_{i}<0)(Y_{i,t}-r_{p}(Z_{i})^{\prime}\beta)^{2}K_{h_{n}}(Z_{i}-\bar{z})$
where $\beta_{+,p,t}=\begin{bmatrix}\mu_{+,t}&\mu_{+,t}^{(1)}&\ldots&\mu_{+,t}^{(p)}/p!\end{bmatrix}^{\prime}$, and $\beta_{-,p,t}=\begin{bmatrix}\mu_{-,t}&\mu_{-,t}^{(1)}&\ldots&\mu_{-,t}^{(p)}/p!\end{bmatrix}^{\prime}$. The estimation procedure is similar to the one described in Section 3. We are examining whether the derivatives of the functions, rather than the levels, are consistent between periods $t=0$ and $t=1$ on each side. To do this, we create a test statistic inspired by the "two-sample" Kolmogorov-Smirnov test:
$D_{+}=\sup(\mathbf{1}_{p+1}-e_{1})\left|\left|\beta_{+,p,1}-\beta_{+,p,0}\right|\right|$
$D_{-}=\sup(\mathbf{1}_{p+1}-e_{1})\left|\left|\beta_{-,p,1}-\beta_{-,p,0}\right|\right|$
The critical values of the test statistic depend on the choice of $p$, so determining the optimal $p$ is on the agenda.
The critical values of the test statistic depend on the choice of $p$. Selecting the right value of $p$ can significantly affect the outcome of the test and the conclusions drawn from the results. Therefore, determining the optimal $p$ is on the agenda.
@@ 5 Monte Carlo Simulations
We now demonstrate the finite-sample properties of the estimator through Monte Carlo simulations and compare the performance of the proposed estimator to that of the local linear RDD estimator proposed by Calonico et al. (2014) and the nonparametric DiD regression estimator proposed by Sant’Anna and Zhao (2020). We consider Data Generating Processes (DGP) based on model 3 from Calonico et al. (2014), with small modifications that will be described.
@@
We conduct our simulation studies in four distinct settings: with time-invariant confounding factors at the threshold, and without any confounding factors, which is the typical scenario for RDDs, for both time-invariant DGPs and time-changing DGPs. For each simulation, we conduct 1000 replications and for each replication, we consider a sample size $n=500$, with $Z_{i}\sim(2\mathcal{B}(2,4)-1)$ where $\mathcal{B}(p_{1},p_{2})$ is a beta distribution with parameters $p_{1}$ and $p_{2}$. We also consider $\varepsilon_{it}\sim N(0,\sigma^{2}_{\varepsilon})$, $\sigma_{\varepsilon}=0.1295$ and the outome generated is $Y_{i,t}=\mu_{it}(Z_{i})+\varepsilon_{i,t}$. The detailed specifications and further functional forms of both models can be found in the Appendix D for added clarity. In both scenarios, we observe that our DiDC estimator performs better than the RDD estimator, with a smaller bias and improved coverage. Additionally, we find that the DiDC estimator has a similar bias and improved coverage when compared to the DiD estimator if the DGPs change over time.
5.1 Identical DGPs over time
@@
In the first simulation, we mimic a scenario where one or more confounding discontinuities are present at the threshold $z_{0}=0$ but the DGPs are the same in both periods. This setting is comparable to that of Grembi et al. (2016), where the treatment of interest was introduced at some point between $t=0$ and $t=1$ and was given to units whose running variable values $Z_{i}$ are above the threshold $z_{0}=0$ however there were other pre-existing treatments determined by the same threshold $z_{0}=0$ on the same running variable.
@@
The second model we consider is a scenario with identical DGPs over time, the only distinction being that in period $t=1$, there is an effect of treatment $\tau$ for units with $Z_{i}\geq 0$. Importantly, there are no other sources of discontinuity in the outcome at the threshold $z_{0}=0$, making this what would be considered an ideal scenario for estimating the effect of the treatment with a regular RDD estimation.
Figure 1: DGPs for models 1 and 2
@@
We present the results of these simulation studies in Table 1 and Table 2. We estimate the DiD along with the RDD and DiD and compare the average bias, median bias, root-mean-squared errors, 95% coverage probability, and 95% confidence interval length for each estimator when the treatment effect $\tau$ is equal to 1.
Table 1: Monte Carlo Simulations - Identical DGPs - Yes Confounder
Av. Bias RMSE Coverage CIL Av. SE Bandwiths
@@
RDD (Robust) 1.042 1.042 0 0.315 0.080 0.183 0.328
DiDC (RDD of $\Delta$s) 0.001 0.061 0.928 0.278 0.070 0.215 0.348
DiDC ($\Delta$ of RDDs) 0.001 0.068 0.924 0.301 0.077 0.183 0.328
DiD 0.001 0.017 1 0.559 0.143 - -
OLS of $\Delta$s 0.001 0.023 1 0.144 0.258 - -
Note: Simulations based on 10,000 Monte Carlo experiments with a sample size $n=500$. RDD is the non-bias corrected RD estimator from CCT (2014), DiD is the outcome regression DiD estimator from Sant’Anna and Zhao (2020), DiDC is the estimator proposed in this paper. “Av. Bias”, “RMSE”, “Cover”, “CIL’ and "Av. SE, stand for the average simulated bias, simulated root mean-squared errors, 95% coverage probability, 95% confidence interval length and average standard error, respectively. "Bandwidths" $h_{n}$ and $b_{n}$ report the plug-in bandwidths for point and bias estimation, respectively.
Table 1 shows that there is a significant improvement over the RDD estimation when there is a confounding factor at the threshold. The reason for this is that the RDD method relies solely on post-treatment data and fails to account for pre-existing confounding factors, making it unsuitable for this setting. In comparison to the DiD, our estimator has closely the same bias, but the coverage seems to be more desirable since it is more informative. While the coverage of DiD is 1, meaning it covers the entire interval of the data, the coverage of the DiDC for the 95% interval is very close to 95%, at 92.8%.
[…]
Table 2: Monte Carlo Simulations - Identical DGPs - No Confounder
Av. Bias RMSE Coverage CIL Av. SE Bandwiths
@@
RDD (Robust) 0.044 0.073 0.903 0.314 0.080 0.183 0.328
DiDC (RDD of $\Delta$s) 0.001 0.060 0.924 0.278 0.070 0.215 0.348
DiDC ($\Delta$ of RDDs) 0.001 0.067 0.919 0.301 0.077 0.183 0.328
DiD 0.001 0.017 1 0.559 0.143 - -
OLS of $\Delta$s 0.001 0.023 1 0.144 0.258 - -
Note: Simulations based on 10,000 Monte Carlo experiments with a sample size $n=500$. RDD is the non-bias corrected RD estimator from CCT (2014), DiD is the outcome regression DiD estimator from Sant’Anna and Zhao (2020), DiDC is the estimator proposed in this paper. “Av. Bias”, “RMSE”, “Cover”, “CIL’ and "Av. SE, stand for the average simulated bias, simulated root mean-squared errors, 95% coverage probability, 95% confidence interval length and average standard error, respectively. "Bandwidths" $h_{n}$ and $b_{n}$ report the plug-in bandwidths for point and bias estimation, respectively.
Table 2 reveals that once again our estimator has a similar bias to that of the DiD, but with more informative coverage. However, the key takeaway is that this table provides a valuable comparison between the DiDC and the RDD. In this simulation, as there were no other discontinuities at the threshold besides the treatment assignment probabilities, the scenario would be considered for the traditional use of RDD.
The results show that difference-in-discontinuities estimation has an apparent improvement over the RDD, yielding a smaller bias and better coverage. This is noteworthy as it highlights that even in cases where the RDD would traditionally be regarded as suitable, the differences-in-discontinuity approach can give more desirable results by incorporating more data into the estimation. The results hold when the functional format exhibits little temporal variation, as shown in the next section, with the additional period contributing to more reliable estimations compared to RDD, whenever the bias from the extra period $t=0$ does not surpass the one from period $t=1$.
5.2 Time-varying DGPs
Next, we introduce scenarios where the Data Generating Processes (DGPs) change between periods, contrasting to the prior section where DGPs were identical across time. This setup allows us to assess the performance of the difference-in-discontinuities in scenarios with small temporal DGP variations.
Figure 2: DGPs for models 3 and 4
We again do a version with confounder and a version without confounder, but now we alter the model for the period $t=1$ so that it is a linear model derived from the original. The DGP for time $t=0$ is identical to that of models 1 and 2. Results are shown in Table 3.
[…]
Table 3: Monte Carlo Simulations - Different DGPs
Av. Bias RMSE Coverage CIL Av. SE Bandwiths

### v2

@@ 4.2 Testing the Time-Invariance of Potential Outcomes Functional Form
Therefore, it can be very beneficial for researchers to know if the shapes of the data-generating functions remain stable over time. This knowledge would allow them to determine whether they can use simpler, less biased estimation methods. We propose a simple nonparametric Two-Sample Kolmogorov-Smirnov (KS) test for time-invariance of the functional forms of conditional means. The procedure is an adaptation of the KS test for conditional moment restrictions from Whang (2001).
In order to implement the test, it is required to have data from more than one time period before the implementation of the treatment of interest ($t\in\left\{-1,0,1\right\}$). Using the notation from Appendix A, we write the nonparametric regression model for the outcome in time $t$ as
$Y_{i,t}=\mu_{t}(Z_{i})+\varepsilon_{i,t}$
where $\mu_{t}(Z_{i})=\mathbb{E}\left[Y_{i,t}|Z_{i}\right]$ and $\varepsilon_{i,t}=Y_{i,t}-\mathbb{E}\left[Y_{i,t}|Z_{i}\right]$. Let $\mathcal{S}^{+}(Z)$ and $\mathcal{S}^{-}(X)$ denote, respectively, the support of the running variable above and below the threshold $z_{0}$. We want to test whether $\mathbb{E}\left[\mu_{0}(Z_{i})-\mu_{-1}(Z_{i})|Z_{i}\right]=0$ almost surely for $z\in\mathcal{S}^{+}(Z)$ and whether $\mathbb{E}\left[\mu_{0}(Z_{i})-\mu_{-1}(Z_{i})|Z_{i}\right]=0$ almost surely for $z\in\mathcal{S}^{-}(Z)$. To do so, we assume the conditional mean of $Y_{i,t}$ can be approximated by a $K\times 1$ vector of approximating functions $p^{K}(.)=(p_{1}(.),...,p_{K}(.))^{{}^{\prime}}$. Using this vector, we write the regression model for the outcome in period $t$ above and below the threshold, respectively, as
$Y_{i,t}=p^{K}(Z_{i})^{{}^{\prime}}\gamma^{+}_{t}+u_{i,t}^{+},$
$Y_{i,t}=p^{K}(Z_{i})^{{}^{\prime}}\gamma^{-}_{t}+u_{i,t}^{-}$
Under the null hypothesis of time-invariant conditional means, we have
$\mathbb{E}\left[\left(p^{K}(Z_{i})^{{}^{\prime}}\gamma^{+}_{0}-p^{K}(Z_{i})^{{}^{\prime}}\gamma^{+}_{-1}\right)\mathbf{1}\left\{Z_{i}\leq z\right\}\right]=0,\forall z\in\mathcal{S}^{+}(Z),$
$\mathbb{E}\left[\left(p^{K}(Z_{i})^{{}^{\prime}}\gamma^{-}_{0}-p^{K}(Z_{i})^{{}^{\prime}}\gamma^{-}_{-1}\right)\mathbf{1}\left\{Z_{i}\leq z\right\}\right]=0,\forall z\in\mathcal{S}^{-}(Z)$
which motivates the following KS type test statistics:
$KS_{+}=\sup_{z\in\mathcal{S}^{+}(z)}\frac{1}{\sqrt{n}^{+}}\sum_{i:Z_{i}\geq z_{0}}\left|\left(p^{K}(Z_{i})^{{}^{\prime}}\widehat{\gamma}^{+}_{0}-p^{K}(Z_{i})^{{}^{\prime}}\widehat{\gamma}^{+}_{-1}\right)\mathbf{1}\left\{Z_{i}\leq z\right\}\right|$
$KS_{-}=\sup_{z\in\mathcal{S}^{-}(z)}\frac{1}{\sqrt{n}^{-}}\sum_{i:Z_{i}\geq z_{0}}\left|\left(p^{K}(Z_{i})^{{}^{\prime}}\widehat{\gamma}^{-}_{0}-p^{K}(Z_{i})^{{}^{\prime}}\widehat{\gamma}^{-}_{-1}\right)\mathbf{1}\left\{Z_{i}\leq z\right\}\right|$
where $n^{+}=\sum_{i=1}^{n}\mathbf{1}\left\{Z_{i}\geq z_{0}\right\}$, $n^{-}=\sum_{i=1}^{n}\mathbf{1}\left\{Z_{i}<z_{0}\right\}$, $\widehat{\gamma}^{+}_{t}$ is the vector of least square estimates of the regression of $Y_{i,t}$ on $p^{K}(Z_{i})^{{}^{\prime}}$ for units above the threshold and $\widehat{\gamma}^{-}_{t}$ is the vector of least square estimates of the regression of $Y_{i,t}$ on $p^{K}(Z_{i})^{{}^{\prime}}$ for units below the threshold.
Confidence intervals and p-values for the test statistics can be obtained using a recentered bootstrap (Hall and Horowitz, 1996) in which the null hypothesis is imposed for the bootstrapped distribution. In Appendix F, we provide the conditions under which the bootstrap is consistent, and the test statistics are powerful against local alternatives.
5 Partial Identification and Sensitivity Analysis
5.1 Partial Identification under Bounded Variation Assumptions
In this section, we consider the partial identification of treatment effects when Assumption 4 is violated. In many settings, assuming that confounding effects remain constant over time is implausible. If that is the case, then the effect of the treatment of interest is not point identified.
We replace the time-invariance assumption with bounded variations assumptions, in the spirit of Manski and Pepper (2018), and derive identified sets for the causal effect of interest as a function of the sensitivity parameters that bound variations in potential outcomes across time. Formally, we invoke the following assumption:
Assumption 6 (Bounded Variations).
Let $c_{1}$ and $c_{2}$ be scalars. We assume that
$\left|\mathbb{E}\left[Y_{i,1}(1,0)-Y_{i,0}(1,0)|Z_{i}=z\right]\right|\leq c_{1},$
$\left|\mathbb{E}\left[Y_{i,1}(1,0)-Y_{i,1}(0,0)|Z_{i}=z_{0}\right]-\mathbb{E}\left[Y_{i,0}(1,0)-Y_{i,0}(0,0)|Z_{i}=z_{0}\right]\right|\leq c_{2}$
The first inequality in Assumption 6 states that the difference between the mean potential outcome $Y_{i,0}(0,1)$ (observed) and the mean potential outcome $Y_{i,1}(1,0)$ (not observed) at the threshold is no greater than the scalar $c_{1}$. Thus, the parameter $c_{1}$ can be interpreted as a time-trend in the evolution of the potential outcome associated to receiving the confounder, but not the treatment of interest.
The second inequality states that the difference between the mean confounding effect in period 0 (point identified under Assumptions 1 and 4), and the mean confounding effect in period 1 (not observed nor identified) at the threshold is no greater than the scalar $c_{2}$. In that sense, $c_{2}$ can be interpreted as a time-trend in confounding effects. In the next lemma, we derive the identified set for $\tau_{c}$:
Lemma 6.
Under Assumptions 1,2 and 5, $\tau_{c}\in\left[\tau_{c}^{LB},\tau_{c}^{UB}\right]$, where
$\tau_{c}^{LB}=\max\left\{\Delta Y^{+}-c_{1},(\Delta Y^{+}-\Delta Y^{-})-c_{2}\right\}$
$\tau_{c}^{UB}=\min\left\{\Delta Y^{+}+c_{1},(\Delta Y^{+}-\Delta Y^{-})+c_{2}\right\}$
Lemma 6 shows that average treatment effect for confounded individuals at the cutoff is partially identified as a function of the sensitivity parameters $c_{1}$ and $c_{2}$. It is straightforward to identify breakdown values for this identified set. That is, the largest values of $c_{1}$ and $c_{2}$ for which a particular conclusion holds. For instance, researchers might be interested in the largest values of bounded variation under which one can conclude that the causal effect of interest is positive. If the lower bound is above 0 for large values of $c_{1}$ and $c_{2}$, then one can conclude that the qualitative conclusions of the empirical exercise are robust to violations of the identifying assumptions.
Estimation of the bounds for $\tau$ is straightforward. We a nonparametric estimator in which the local polynomial estimates described in Section 3 are plugged in the max and min operators. The estimators for the lower and the upper bound are, respectively,
$\widehat{\tau}_{c}^{LB}=\max\left\{\Delta\hat{\mu}^{(\nu)}_{+,p}(h_{n},b_{n})-c_{1},(\Delta\hat{\mu}^{(\nu)}_{+,p}(h_{n},b_{n})-\Delta\hat{\mu}^{(\nu)}_{-,p}(h_{n},b_{n}))-c_{2}\right\}$
$\widehat{\tau}_{c}^{UB}=\min\left\{\Delta\hat{\mu}^{(\nu)}_{+,p}(h_{n},b_{n})+c_{1},(\Delta\hat{\mu}^{(\nu)}_{+,p}(h_{n},b_{n})-\Delta\hat{\mu}^{(\nu)}_{-,p}(h_{n},b_{n}))+c_{2}\right\}$
We use the Delta method for Hadamard directionally differentiable mappings Fang and Santos (2018) to derive the asymptotic properties of the plug-in estimators for the bounds. The next theorem shows that the estimators of the bounds converge to a non-Gaussian limiting process at the $\sqrt{n\min\left\{h_{n},b_{n}\right\}}$ rate:
Theorem 2.
Suppose Assumptions 1, 3 and 6 hold. Furthermore, assume that $c_{1},c_{2}\in\mathcal{C}$ for some finite grid $\mathcal{C}$. Then,
$\sqrt{n\min\left\{h_{n},b_{n}\right\}}\begin{pmatrix}\widehat{\tau}_{c}^{LB}-\tau_{c}^{LB}\\ \widehat{\tau}_{c}^{UB}-\tau_{c}^{UB}\end{pmatrix}\rightarrow\textbf{Z}(y,z,c_{1},c_{2})$
a tight random element of $l^{\infty}\left(\mathcal{S}(Y)\times\mathcal{S}(Z)\times\mathcal{C},\mathbb{R}^{2}\right)$.
Inference for the bounds estimates is particularly challenging. Using the limiting process to which the estimates converge to obtain analytical asymptotic confidence bands is difficult. An alternative would be a bootstrap. Although the Delta Method is valid for Hadamard directionally differentiable functions, the standard nonparametric bootstrap is not. Instead, we recommend researchers to use the bootstrap procedure proposed in Section3.2.1 in Fang and Santos (2018) [^2] Note, however, that the mapping is fully differentiable if the quantities within the min and max operators are different, in which case the standard bootstrap is valid.
5.2 Partial Identification under Modularity Assumptions
An alternative approach to partial identification in the Diff-in-Disc setting is to partially identify treatment effects by exploiting assumptions regarding the interaction between the treatment of interest and the confounding policy. The first step is to assume that potential outcomes are bounded up to known values:
Assumption 7 (Bounded Outcomes).
For all $i\in\left\{1,...,n\right\}$, $t\in\left\{0,1\right\}$ and $(d_{0},d_{1})\in\left\{0,1\right\}^{2}$,
$\infty<y^{min}\leq Y_{i,t}(d_{0},d_{1})\leq y^{max}<\infty$
In some settings the values $y^{min}$ and $y^{max}$ readily justified by the nature of the outcome, and in others these values are sensitivity parameters that reflect the beliefs about the smallest and largest possible values of potential outcomes (Kline and Masten, 2025).
Assumption 7 motivates the worst-case bounds (Manski, 1989). Without further assumptions, it implies that the treatment effects lie in the interval $\left[y^{min}-y^{max},y^{max}-y^{min}\right]$. Such bounds are usually uninformative as the bound covers zero, and hence the sign of the treatment effect is not identified.
Theoretical knowledge of the applied setting can be used to tighten the bounds. For instance, researchers can draw insights from economic theory to assume that the treatment of interest and the confounding policy are complementary:
Assumption 8 (Complementarity).
For all $i\in\left\{1,...,n\right\}$ and $t\in\left\{0,1\right\}$,
$Y_{i,1}(1,0)+Y_{i,1}(0,1)\leq Y_{i,1}(1,1)+Y_{i,1}(0,0)$
Assumption 8 is the supermodularity assumption from Twinam (2017) for the case of two binary treatments. It formalizes the notion that the magnitude of the causal effect of interest increases with the confounding policy. Mapping the assumption to our empirical setting, one can assume, for instance, that the effect of relaxing fiscal rules on public finance outcomes is greater for municipalities with less-skilled incumbents. The proposition below derives the sharp identified sets for $\tau_{c}$ and $\tau_{uc}$ under complementarity:
Lemma 7.
Suppose Assumptions 1, 3, 7 and 8 hold. Then,
$\tau_{c}\in\left[y^{min}-Y_{1}^{-},y^{max}-y^{min}\right]$
and
$\tau_{uc}\in\left[y^{min}-y^{max},Y_{1}^{+}-y^{min}\right]$
[…]
Alternatively, researchers can assume that the treatment of interest and the confounder are substitutes:
Assumption 9 (Substitutability).
For all $i\in\left\{1,...,n\right\}$ and $t\in\left\{0,1\right\}$,
$Y_{i,t}(1,0)+Y_{i,t}(0,1)\geq Y_{i,t}(1,1)+Y_{i,t}(0,0)$
Assumption 9 is the submodularity assumption from Twinam (2017) for the case of two binary treatments. It states that the treatment effect of interest decreases in the confounding policy. The proposition below derives the identified sets for $\tau_{c}$ and $\tau_{uc}$ under substitutability:
Lemma 8.
Suppose Assumptions 1, 3, 7 and 9 hold. Then,
$\tau_{c}\in\left[y^{min}-y^{max},y^{max}-Y_{1}^{-}\right]$
and
$\tau_{uc}\in\left[Y_{1}^{+}-y^{max},y^{max}-y^{min}\right]$
5.3 Partial Identification Combining Bounded Variation and Modularity Assumptions
So far in this section, we have described two approaches to partially identify the target parameters in the DiDC setting under different sets of assumptions.
The bounded variation approach partially identifies $\tau_{c}$ as a function of a measure of deviation from Assumption 4. The procedure has several desirable features. It allows for researchers to partially identify the parameters under different values for the deviation, which also allows researchers to assess the robustness of empirical findings. However, it does not allow for the partial identification of $\tau_{uc}$, not at least without further assumptions regarding the homogeneity of treatment effects.
The modularity approach, on the other hand, allows for the partial identification of $\tau_{c}$ ad $\tau_{uc}$, by invoking assumptions on the interaction of the confounding policy and the treatment of interest which can be rooted in economic theory or background knowledge of the empirical setting. However, the procedure does not allow researchers to exploit the temporal structure of the data. Nevertheless, it is (particularly) useful for partial identification in RD settings in which there is a confounding policy at the threshold, but no data from periods before the implementation of the treatment of interest.
In the lemmas below, we show that combining bounded variation and modularity assumptions can strengthen the bounds on $\tau_{uc}$ and $\tau_{c}$:
Lemma 9.
Suppose Assumptions 1, 3 and 6-8 hold. Then, $\tau_{c}\in\left[\tau_{c}^{LB},\tau_{c}^{UB}\right]$ and $\tau_{uc}\in\left[\tau_{uc}^{LB},\tau_{uc}^{UB}\right]$, where
$\tau_{c}^{LB}=\max\left\{\Delta Y^{+}-c_{1},(\Delta Y^{+}-\Delta Y^{-})-c_{2},y^{min}-Y_{1}^{-}\right\}$
$\tau_{c}^{UB}=\min\left\{\Delta Y^{+}+c_{1},(\Delta Y^{+}-\Delta Y^{-})+c_{2},y^{max}-y^{min}\right\}$
and
$\tau_{uc}^{LB}=y^{min}-y^{max}$
$\tau_{uc}^{UB}=\min\left\{\Delta Y^{+}+c_{1},Y_{1}^{+}-y^{min}\right\}$
Lemma 10.
Suppose Assumptions 1, 3 and 6, 7 and 9 hold. Then, $\tau_{c}\in\left[\tau_{c}^{LB},\tau_{c}^{UB}\right]$ and $\tau_{uc}\in\left[\tau_{uc}^{LB},\tau_{uc}^{UB}\right]$, where
$\tau_{c}^{LB}=\max\left\{\Delta Y^{+}-c_{1},(\Delta Y^{+}-\Delta Y^{-})-c_{2},y^{min}-y^{max}\right\}$
$\tau_{c}^{UB}=\min\left\{\Delta Y^{+}+c_{1},(\Delta Y^{+}-\Delta Y^{-})+c_{2},y^{max}-Y_{1}^{-}\right\}$
and
$\tau_{uc}^{LB}=\max\left\{\Delta Y^{+}-c_{1},Y_{1}^{+}-y^{max}\right\}$
$\tau_{uc}^{UB}=y^{max}-y^{min}$
Lemmas 9 and 10 derive the identified sets for $\tau_{c}$ and $\tau_{uc}$ when bounded variation, bounded outcomes and modularity assumptions are combined. When it comes to $\tau_{c}$, the combination of assumptions strengthens the lower and the upper bound under both modularity assumptions. When it comes to $\tau_{uc}$, however, only one of the bounds is strengthened by the combination of assumptions, whereas the other bounds remains the worst-case bound. When it comes to estimation and inference, the bootstrap procedure outlined in Theorem 2 remains valid for bounds under alternative sets of assumptions as well.
@@ 6 Monte Carlo Simulations
We now analyze the finite-sample properties of the Diff-in-Disc estimator through Monte Carlo simulations and compare the performance of the proposed estimator to that of the local linear RDD estimator proposed by Calonico et al. (2014b), the nonparametric DiD regression estimator proposed by Sant’Anna and Zhao (2020), and the Two-Way Fixed Effects (TWFE) estimator. We consider Data Generating Processes (DGP) based on model 3 from Calonico et al. (2014b), with small modifications that will be described.
@@
We conduct our simulation studies in four distinct settings: with time-invariant confounding factors at the threshold and without any confounding factors, which is the typical scenario for RDDs, for both time-invariant and time-varying functional forms of the conditional means of potential outcomes. For each simulation, we conduct 1000 replications, and for each replication, we consider a sample size $n=1000$, with $Z_{i}\sim(2\mathcal{B}(2,4)-1)$ where $\mathcal{B}(p_{1},p_{2})$ is a beta distribution with parameters $p_{1}$ and $p_{2}$. We also consider $\varepsilon_{it}\sim N(0,\sigma^{2}_{\varepsilon})$, $\sigma_{\varepsilon}=0.1295$ and the outcome generated is $Y_{i,t}=\mu_{it}(Z_{i})+\varepsilon_{i,t}$. The detailed specifications and additional functional forms of both models are provided in Appendix D for clarity. In both scenarios, we observe that our DiDC estimator performs better than the RDD estimator, with a smaller bias and improved coverage. Additionally, we find that the DiDC estimator has smaller bias and better coverage than the DiD estimator when the DGPs change over time.
6.1 Identical Functional Forms over Time
@@
In the first simulation, we mimic a scenario where one or more confounding discontinuities are present at the threshold $z_{0}=0$, but the functional forms for conditional means of potential outcomes are the same in both periods. This setting is comparable to that of Grembi et al. (2016), where the treatment of interest was introduced at some point between $t=0$ and $t=1$ and was given to units whose running variable values $Z_{i}$ are above the threshold $z_{0}=0$ however there were other pre-existing treatments determined by the same threshold $z_{0}=0$ on the same running variable.
@@
The second model we consider is a scenario with identical functional forms over time, the only distinction being that in period $t=1$, there is an effect of treatment $\tau$ for units with $Z_{i}\geq 0$. Importantly, there are no other sources of discontinuity in the outcome at the threshold $z_{0}=0$, making this an ideal scenario for estimating the effect of the treatment using a regular RDD.
@@
We present the results of these simulation studies in Tables 1 and 2. We estimate the Diff-in-Disc along with the RDD, DiD and TWFE, and compare the average bias, median bias, root-mean-squared errors, 95% coverage probability, and the length of the 95% confidence interval for each estimator when the treatment effect $\tau$ is equal to 0.
Table 1: Identical Functional Forms with Confounding
AV.Bias RMSE Coverage CIL Bandwidths
@@
RDD (Robust) 1.043 1.043 0 0.175 0.169 0.318
Diff-in-Disc -0.002 0.047 0.944 0.228 0.225 0.360
NP-DiD 0.001 0.019 0.953 0.393 - -
TWFE 0.001 0.011 1 0.397 - -
Note: Simulations based on 10,000 Monte Carlo experiments with a sample size $n=1000$. RDD is the non-bias corrected RD estimator from CCT (2014), NP-DiD is the outcome regression DiD estimator from Sant’Anna and Zhao (2020), DiDC is the estimator proposed in this paper. “Av. Bias”, “RMSE”, “Cover”, “CIL’ stand for the average simulated bias, simulated root mean-squared errors, 95% coverage probability and 95% confidence interval length, respectively. "Bandwidths" $h_{n}$ and $b_{n}$ report the plug-in bandwidths for point and bias estimation, respectively.
Table 1 shows a significant improvement over the RD estimator when a confounding factor is present at the threshold. The result is not surprising, as the presence of the confounder violates the continuity assumption that underlies the validity of cross-sectional RD. Note also that the mean size of the bandwidths, both for the estimator and for the bias-correction, is greater for the Diff-in-Disc estimator than the RD estimator. In terms of bias, both the nonparametric DiD and the TWFE estimators exhibit desirable finite-sample properties. However, when it comes to coverage of the confidence interval, we find that those from the TWFE estimator are severely biased.
Table 2: Identical Functional Forms without Confounding
AV.Bias RMSE Coverage CIL Bandwidths
@@
RDD (Robust) 0.043 0.051 0.861 0.176 0.169 0.318
Diff-in-Disc -0.003 0.047 0.944 0.228 0.223 0.360
NP-DiD 0.001 0.019 0.955 0.391 - -
TWFE 0.002 0.012 1 0.397 - -
Note: Simulations based on 10,000 Monte Carlo experiments with a sample size $n=1000$. RDD is the non-bias corrected RD estimator from CCT (2014), NP-DiD is the outcome regression DiD estimator from Sant’Anna and Zhao (2020), DiDC is the estimator proposed in this paper. “Av. Bias”, “RMSE”, “Cover”, “CIL’ stand for the average simulated bias, simulated root mean-squared errors, 95% coverage probability and 95% confidence interval length, respectively. "Bandwidths" $h_{n}$ and $b_{n}$ report the plug-in bandwidths for point and bias estimation, respectively.
Table 2 shows the results for the case where there is no confounding policy at the threshold. In that case, the cross-sectional RD is valid, as evidenced by the small finite-sample bias in the simulation. Once again, the Diff-in-Disc estimator has smaller finite-sample bias and larger bandwidths, which illustrates the point in Section 3 that, even in settings where the standard RD is valid, there might be gains in using the Diff-in-Disc approach.
The results show that Difference-in-Discontinuites estimator has an apparent improvement over the standard RD, yielding a smaller bias and better coverage. This is noteworthy as it highlights that even in cases where the RDD would traditionally be regarded as suitable, the differences-in-discontinuity approach can give more desirable results by incorporating more data into the estimation. The results hold when the functional format exhibits little temporal variation, as shown in the next section, with the additional period contributing to more reliable estimates than the RD whenever the bias from the extra period $t=0$ does not exceed that from period $t=1$. In Appendix D, we show that the results for the simulations are robust even in the case where the variance of potential outcomes vary over time.
6.2 Time-varying Functional Forms
Next, we introduce scenarios where the functional forms for conditional means change between periods, contrasting to the prior section where functional forms were identical across time. This setup allows us to analyze how changes in the conditional mean of potential outcomes affect the performance of the considered estimators.
Again, we consider simulations with and without confounders at the threshold, but now we alter the model for $t=1$ to be a linear model derived from the original. The functional form for time period $t=0$ is identical to that of models 1 and 2. Results are shown in Tables 3 and 4.
Table 3: Time-Varying Functional Forms with Confounding
AV.Bias RMSE Coverage CIL Bandwidths